PLEASE HELP ME!!
how do you write the equation of a parabola in the standard form y=ax^2+bx+c, given 2 points and the x-intercept??
point 1: (5,30)
point 2: (15,30)
x-intercept: (20,0)

Answer by t0hierry(194)

y - 30 = a(x-5)(x-15) satisfies the first two conditions. Next we use the third constraint, which is
when x=20 we have 0 - 30 = a (20-5)(20-15) = 75a
or a = -30/75 = -10/25 = -0.4
I get y = -0.4 x^2 + 8x + 30
Let me know if this helps